Find a unit vector perpendicular to the plane through P(2, 1, -1), Q(-1, 1, 2) and R (1, -1, 2)


1
Expert's answer
2022-05-25T15:32:53-0400
"\overrightarrow{PQ}=(-3,0,3), \overrightarrow{PR}=(-1,-2,3)"

"\overrightarrow{PQ}\times \overrightarrow{PR}=\begin{vmatrix}
 \vec{i} & \vec{j} &\vec{k} \\
 -3 & 0 & 3 \\
 -1 & -2 & 3 \\
\end{vmatrix}"

"=\vec{i}\begin{vmatrix}
 0 & 3 \\
 -2& 3
\end{vmatrix}-\vec{j}\begin{vmatrix}
 -3 & 3 \\
 -1 & 3
\end{vmatrix}+\vec{k}\begin{vmatrix}
 -3 & 0 \\
 -1 & -2
\end{vmatrix}"

"=6\vec{i}+6\vec{j}+6\vec{k}"

"|\overrightarrow{PQ}\times \overrightarrow{PR}|=6\sqrt{3}"

"\vec{u}=(\dfrac{1}{\sqrt{3}},\dfrac{1}{\sqrt{3}},\dfrac{1}{\sqrt{3}})"